OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 70 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 240 miles apart? 4 hrs 6 hrs 2 hrs 5 hrs

OpenStudy (anonymous):

Helppp!

OpenStudy (ganpat):

s = d/ t (70 - 50 ) = 240/ t 20 = 240/ t t = 6 hrs..

OpenStudy (anonymous):

Thank you!